# Statistical Test on Diffusion Model-based Anomaly Detection by Selective Inference

### v3

@@ 3 Statistical Test on Generated Images by Diffusion Models
where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix. In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image is denoted as $\bm{X}_{\rm obs}$.
Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}:\mathbb{R}^{n}\ni\bm{X}\rightarrow\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$. The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error. When identifying anomalous regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise. In this study, we simply used an averaging filter. Let us denote the averaging filter as $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Then, the process of obtaining the (filtered) reconstruction error is written as
$E:\mathbb{R}^{n}\ni\bm{X}\mapsto|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|\in\mathbb{R}^{n},$ (6)
where absolute value is taken pixel-wise. Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $E_{i}(\bm{X})$ for each pixel $i\in[n]$. Specifically, we define the anomalous region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in(0,\infty)$, i.e.,
$\mathcal{M}_{\bm{X}}=\left\{i\in[n]\mid E_{i}(\bm{X})\geq\lambda\right\}.$ (7)
Statistical inference.
In order to quantify the statistical significance of the anomalous regions detected by a diffusion model, we consider the concrete example of two-sample test. Let us define an $n$-dimensional reference input vector,
$\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}=\bm{\mu}^{\mathrm{ref}}+\bm{\varepsilon}^{\mathrm{ref}},\;\;\;\bm{\varepsilon}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma),$ (8)
@@
$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu^{\rm ref}_{i}\\ \text{v.s.}\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu^{\rm ref}_{i},\end{gathered}$ (9)
@@
where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference images in the anomalous regions, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different. A reasonable test statistic for the statistical test in (9) is the difference in mean pixel values between the test image and the reference image in the anomalous region $\mathcal{M}_{\bm{X}}$, i.e.,
$T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X_{i}-\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X^{\mathrm{ref}}_{i}=\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix},$ (10)
@@
where $\bm{\nu}_{\mathcal{M}_{\bm{X}}}\in\mathbb{R}^{2n}$ is the vector that depends on the anomalous region $\mathcal{M}_{\bm{X}}$, defined as
$\bm{\nu}_{\mathcal{M}_{\bm{X}}}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\begin{pmatrix}\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}\\[5.0pt] -\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}\end{pmatrix}\in\mathbb{R}^{2n},$ (11)
where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise. If we do not account for the fact that the anomalous region is detected by a diffusion model, the distribution of the test statistic would be simply given as
$T(\bm{X},\bm{X}^{\mathrm{ref}})\sim\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}),\;\;\;\text{where}\;\;\tilde{\Sigma}=\begin{pmatrix}\Sigma&O_{n}\\ O_{n}&\Sigma\end{pmatrix}.$ (12)
In this case, the $p$-values defined as
$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|>|T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})|\right),$ (13)
would be easily computed by the normality of the test statistic distribution. However, in reality, since the anomalous region is detected by the trained diffusion model, $\bm{\nu}_{\mathcal{M}_{\bm{X}}}$ depends on the data $\bm{X}$, meaning that the sampling distribution of the test statistic is much more complicated. Therefore, if $p_{\mathrm{naive}}$ is used for decision-making, the false detection error cannot be properly controlled.
4 Computing Selective $p$-values
@@
In this section, we introduce selective inference (SI) framework for testing images generated by diffusion models and propose a method to perform valid hypothesis test.
4.1 Conditional Distribution of Test Statistics
Due to the complexity described in the previous section, it is difficult to directly obtain the sampling distribution of $T(\bm{X},\bm{X}^{\mathrm{ref}})$. Then, we consider the sampling distribution of $T(\bm{X},\bm{X}^{\mathrm{ref}})$ conditional on the event that the anomalous region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomalous region $\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$, i.e.,
$T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}\right\}.$ (14)
In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameter that is independent of the test statistic. As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space. The nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
@@ 4.1 Conditional Distribution of Test Statistics
$\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}\bm{\nu}_{\mathcal{M}_{\bm{X}}}^{\top}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}.$ (15)
The $p$-value calculated from this conditional sampling distribution is called a selective $p$-value. Specifically, the selective $p$-value is defined as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|>|T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})|\mid\bm{X}\in\mathcal{X}\right),$ (16)
where $\mathcal{X}$ is the conditional data space defined as
[…]
Due to the conditioning on the nuisance parameter $\mathcal{Q}_{\bm{X}}$, the conditional data space $\mathcal{X}$ can be rewritten as
$\mathcal{X}=\left\{\begin{pmatrix}\bm{X}(z)\\ \bm{X}^{\mathrm{ref}}(z)\end{pmatrix}\in\mathbb{R}^{2n}\middle|\begin{pmatrix}\bm{X}(z)\\ \bm{X}^{\mathrm{ref}}(z)\end{pmatrix}=\bm{a}+\bm{b}z,z\in\mathcal{Z}\right\},$ (18)
where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$, and $\bm{c}_{1:n}$ represents a vector composed of the first $n$ elements of the vector $\bm{c}$. The vectors $\bm{a},\bm{b}\in\mathbb{R}^{2n}$ are defined as
@@
$\bm{a}=\mathcal{Q}_{\bm{X}_{\mathrm{obs}}},\;\bm{b}=\frac{\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}},$ (19)
and the region $\mathcal{Z}$ is defined as
$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}\right\}.$ (20)
Let us consider a random variable $Z\in\mathbb{R}$ and its observation $z_{\mathrm{obs}}\in\mathbb{R}$ so that they satisfy $\bm{X}=\bm{a}_{1:n}+\bm{b}_{1:n}Z$ and $\bm{X}_{\mathrm{obs}}=\bm{a}_{1:n}+\bm{b}_{1:n}z_{\mathrm{obs}}$. Then, the selective $p$-value in (16) is re-written as
@@
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right).$ (21)
Under the null hypothesis $\mathrm{H}_{0}$, the distribution of the unconditional variable $Z$ is $\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}})$. Consequently, given $Z\in\mathcal{Z}$, the conditional random variable $Z$ adheres to a truncated Gaussian distribution. Once the truncated region $\mathcal{Z}$ is identified, computing the selective $p$-value in (21) becomes straightforward. Therefore, the remaining task is the identification of $\mathcal{Z}$.
4.2 Over-Conditioning
To compute the truncated region $\mathcal{Z}$, we employ a divide and conquer approach. It is difficult to directly identify the truncated region $\mathcal{Z}$ due to the complexity of the computational algorithm of the diffusion model. The basic idea of this approach is to decompose the data space $\mathcal{X}$ into a set of polyhedra by considering additional conditioning, which we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022). It is easy to understand that a polyhedron in the $n$-dimensional data space ${\mathcal{X}}$ corresponds to an interval in the one-dimensional space ${\mathcal{Z}}$. Therefore, we can sequentially examine intervals in the one-dimensional space and check whether the same hypothesis (anomalous region) as the observed one is selected. In this study, we show that the filtered reconstruction error $E(\bm{X})$ can be expressed as a piecewise-linear function of $\bm{X}$. By exploiting this, we identify a over-conditioned interval $\mathcal{Z}^{\mathrm{oc}}\subset\mathcal{Z}$.
Identification of $\mathcal{Z}^{\mathrm{oc}}$.
Let us write a polyhedron $\mathcal{P}$ composed of piecewise-linear functions as
$\mathcal{P}_{k}=\left\{\bm{\Delta}_{k}\bm{X}\leq\bm{\delta}_{k}\right\},k\in[K],$ (22)
where $\bm{\Delta}_{k}$ and $\bm{\delta}_{k}$ for $k\in[K]$ are the coefficient matrix and the constant vector with appropriate dimensions of the $k$-th piecewise-linear function, respectively. Then, a piecewise-linear function $\mathcal{A}(\bm{X})$ is written in the following form:
$\mathcal{A}(\bm{X})=\begin{cases}\bm{\Psi}_{1}\bm{X}+\bm{\psi}_{1}&\text{if }\bm{X}\in\mathcal{P}_{1},\\ \bm{\Psi}_{2}\bm{X}+\bm{\psi}_{2}&\text{if }\bm{X}\in\mathcal{P}_{2},\\ \hskip 28.45274pt\vdots\\ \bm{\Psi}_{K}\bm{X}+\bm{\psi}_{K}&\text{if }\bm{X}\in\mathcal{P}_{K},\end{cases}$ (23)
where $\bm{\Psi}_{k}$ and $\bm{\psi}_{k}$ for $k\in[K]$ are the coefficient matrix and the constant vector with appropriate dimensions for the $k$-th polyhedron, respectively. Using the notation in (18), since the input image $\bm{X}(z)$ is restricted on a one-dimensional line, each component of the output of $\mathcal{A}$ is written as
$\begin{gathered}\mathcal{A}_{i}(\bm{X}(z))=\begin{cases}\kappa^{\mathcal{A}_{i}}_{1}z+\rho^{\mathcal{A}_{i}}_{1}&\text{if }[L^{\mathcal{A}_{i}}_{1},U^{\mathcal{A}_{i}}_{1}],\\ \kappa^{\mathcal{A}_{i}}_{2}z+\rho^{\mathcal{A}_{i}}_{2}&\text{if }[L^{\mathcal{A}_{i}}_{2},U^{\mathcal{A}_{i}}_{2}],\\ \hskip 28.45274pt\vdots\\ \kappa^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}z+\rho^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}&\text{if }[L^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})},U^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}],\end{cases}\end{gathered}$ (24)
where $K(\mathcal{A}_{i})$ is the number of linear pieces of the piecewise-linear function, and $\kappa^{\mathcal{A}_{i}}_{k}\in\mathbb{R}$ and $\rho^{\mathcal{A}_{i}}_{k}\in\mathbb{R}$ for $k\in[K(\mathcal{A}_{i})]$ are the coefficient and the constant of the $k$-th polyhedron, respectively. For each $i\in[n]$, there exists $k\in[K(\mathcal{A}_{i})]$ such that $z\in[L^{\mathcal{A}_{i}}_{k},U^{\mathcal{A}_{i}}_{k}]$, then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
$\begin{gathered}\coloneqq\begin{cases}\left[\max\left(L^{\mathcal{A}_{i}}_{k},\left((\lambda-\rho^{\mathcal{A}_{i}}_{k})/\kappa^{\mathcal{A}_{i}}_{k}\right)\right),U^{\mathcal{A}_{i}}_{k}\right]&\text{if }\kappa^{\mathcal{A}_{i}}_{k}>0,\\ \left[L^{\mathcal{A}_{i}}_{k},\min\left(U^{\mathcal{A}_{i}}_{k},\left((\lambda-\rho^{\mathcal{A}_{i}}_{k})/\kappa^{\mathcal{A}_{i}}_{k}\right)\right)\right]&\text{if }\kappa^{\mathcal{A}_{i}}_{k}<0.\end{cases}\end{gathered}$ (25)
We denote the over-conditioned interval as
$\mathcal{Z}^{\text{oc}}(\bm{a}+\bm{b}z)=\bigcap_{i\in[n]}\left[L^{i}_{z},U^{i}_{z}\right].$ (26)
Piecewise linearity of diffusion models.
We now show that the diffusion model mapping $\mathcal{D}$ and then filtered reconstruction error $E$ can be expressed as a piecewise-linear function of $\bm{X}$. To show this, we see that both the forward process and reverse process of the diffusion model are piecewise-linear functions as long as we employ a class of U-Net described below. It is easy to see the piecewise-linearity of the forward process as long as we fix the random seed for $\epsilon_{t}$. To make the reverse process a piecewise-linear function, we employ a U-Net architecture composed of piecewise-linear components such as ReLU activation function and average pooling. Then, $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$ is represented as a piecewise-linear function of $\mathbf{x}_{t}$. Moreover, since $f^{(t)}_{\theta}(\mathbf{x}_{t})$ in (3) is a composite function of $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$, it is also a piecewise-linear function. By combining them together, we see that $\mathbf{x}_{t-1}$ is written as a piecewise-linear function of $\mathbf{x}_{t}$. Therefore, the entire reconstruction process is a piecewise-linear function since it just repeats the above operation multiple times (see Algorithm 1). As a result, the entire mapping ${\mathcal{D}}(\mathbf{X})$ of the diffusion model is a piecewise-linear function of the input image $\bm{X}$. Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=E(\bm{X}))$ is piecewise-linear. By exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\text{oc}}$ can be computed.
4.3 Identification of $\mathcal{Z}$ by Parametric Programming

### v4

@@ 3 Statistical test for diffusion-based anomalous localization
where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix. In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image vector is denoted as $\bm{x}$.
Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}\colon\mathbb{R}^{n}\ni\bm{X}\to\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$. The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error. When identifying anomalous regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise. In this study, we simply used an averaging filter. Let us denote the averaging filter as $\mathcal{F}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$. Then, the process of obtaining the (filtered) reconstruction error is written as
$\mathcal{E}\colon\mathbb{R}^{n}\ni\bm{X}\mapsto|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|\in\mathbb{R}^{n},$ (7)
where absolute value is taken pixel-wise. Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $\mathcal{E}_{i}(\bm{X})$ for each pixel $i\in[n]$. Specifically, we define the anomalous region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in\mathbb{R}^{+}$, i.e.,
$\mathcal{M}_{\bm{X}}=\left\{i\in[n]\mid\mathcal{E}_{i}(\bm{X})\geq\lambda\right\}.$ (8)
Statistical test for diffusion models
In order to quantify the statistical significance of the anomalous regions detected by using the diffusion model, we consider the concrete example of two-sample test. Note that our method can be extended to other statistical tests using various statistics. In the two-sample test, we compare the test image and the randomly chosen reference image in the anomalous region. Let us define an $n$-dimensional reference input vector,
$\begin{gathered}\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}=\bm{\mu}^{\mathrm{ref}}+\bm{\varepsilon}^{\mathrm{ref}},\;\;\text{with}\;\;\bm{\varepsilon}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma),\end{gathered}$ (9)
@@
$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i}\\ ~{}~{}~{}\text{v.s.}~{}~{}~{}\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i},\end{gathered}$ (10)
@@
where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference images in the observed anomalous region $\mathcal{M}_{\bm{x}}$, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different. A reasonable test statistic for the statistical test in (10) is the difference in mean pixel values between the test image and the reference image in the anomalous region $\mathcal{M}_{\bm{x}}$, i.e.,
$T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}X_{i}-\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}X^{\mathrm{ref}}_{i}=\bm{\nu}^{\top}\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix},$ (11)
@@
where $\bm{\nu}\in\mathbb{R}^{2n}$ denotes a vector that depends on the anomalous region $\mathcal{M}_{\bm{x}}$, and hence on $\bm{x}$ itself, and is defined as
$\bm{\nu}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\begin{pmatrix}\bm{1}^{n}_{\mathcal{M}_{\bm{x}}}\\[5.0pt] -\bm{1}^{n}_{\mathcal{M}_{\bm{x}}}\end{pmatrix}\in\mathbb{R}^{2n},$ (12)
where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise. In this case, the $p$-value called naive $p$-value, and defined as
$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\right).$ (13)
If we can identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$, we can compute a valid $p$-value that control the false positive detection rate (i.e., the type I error rate).
4 Selective inference for diffusion-based anomaly localization
@@
In this section, we introduce selective inference (SI) framework for testing the anomalous regions detected by a diffusion model and propose a method to perform valid hypotheses tests.
4.1 Computing valid $p$-values via selective inference
Complexity of sampling distribution
As mentioned in §3, we need to identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$ to compute the $p$-values. If we ignore the fact that the anomalous region is identified by a diffusion model, the test statistic in (11) is simply a linear transformation of the Gaussian random vectors $\bm{X}$ and $\bm{X}^{\mathrm{ref}}$, and hence itself follows a Gaussian distribution:
@@ 4.1 Computing valid $p$-values via selective inference
$\begin{gathered}T(\bm{X},\bm{X}^{\mathrm{ref}})\sim\mathcal{N}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}),\;\;\text{where}\;\;\tilde{\Sigma}=\begin{pmatrix}\Sigma&O_{n}\\ O_{n}&\Sigma\end{pmatrix}.\end{gathered}$ (14)
However, as mentioned in §3, in reality the dependence of $\bm{\nu}$ on $\bm{x}$ via the diffusion model is more intricate, making the sampling distribution intractably complex. Consequently, obtaining this sampling distribution directly is challenging.
Selective $p$-value via conditional sampling distribution
Then, we consider the sampling distribution of the test statistic conditional on the event that the anomalous region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomalous region $\mathcal{M}_{\bm{x}}$, i.e.,
$T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}}\right\}.$ (15)
In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameter that is independent of the test statistic. As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space. The sufficient statistics of the nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
@@
$\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}\bm{\nu}^{\top}}{\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix},$ (16)
where $I_{2n}$ is the identity matrix of size $2n$. By additional conditioning on the nuisance parameter, the selective $p$-value is defined as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right).$ (17)
The following theorem establishes that the selective $p$-value is a valid $p$-value for controlling the false positive detection rate for any significance level $\alpha\in(0,1)$.
Theorem 1.
The selective $p$-value in (17) is valid for controlling the false positive detection rate, i.e,
@@
$\mathbb{P}_{\mathrm{H}_{0}}\left(p_{\mathrm{selective}}\leq\alpha\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right)=\alpha,\;\forall\alpha\in(0,1).$ (18)
Then, the selective $p$-value satisfies the following condition:
$\mathbb{P}_{\mathrm{H}_{0}}(p_{\mathrm{selective}}\leq\alpha)=\alpha,\;\forall\alpha\in(0,1).$ (19)
The proof of Theorem 1 is given in Appendix A.1. The following theorem tells that the selective $p$-value can be analytically derived from the conditional sampling distribution, which follows a truncated Gaussian distribution.
Theorem 2.
Consider the truncation intervals defined as
$\mathcal{Z}=\left\{z\in\mathbb{R}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\},$ (20)
where $X(z)$ are defined as
$\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z,\;\;\text{where}\;\;\bm{a}=\mathcal{Q}_{\bm{x}},\;\bm{b}=\frac{\tilde{\Sigma}\bm{\nu}}{\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}},$ (21)
and $\bm{a}_{1:n}$ and $\bm{b}_{1:n}$ denote the first $n$ elements of $\bm{a},\bm{b}\in\mathbb{R}^{2n}$ , respectively. Then, the selective $p$-value in (17) can be rewritten as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}\right).$ (22)
The conditional sampling distribution of the test statistic $T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))\mid\{Z\in\mathcal{Z}\}$ follows a truncated Gaussian distribution $\mathcal{TN}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu})$.
The proof of the Theorem 2 is given in Appendix A.2. Once the truncation intervals $\mathcal{Z}$ are identified, computing the selective $p$-value in (22) becomes straightforward. Therefore, the remaining task is the identification of $\mathcal{Z}$.
4.2 Identification of subinterval $\mathcal{Z}^{\mathrm{sub}}$ in $\mathcal{Z}$
To tackle the identification of the truncation invervals $\mathcal{Z}$, we employ a divide-and-conquer strategy. Directly characterizing $\mathcal{Z}$ is challenging due to the complexity of the diffusion model’s computation. Our method decomposes the $n$-dimensional data space into a collection of polyhedra by imposing additional conditioning, a process we refer to as over-conditioning (OC) Duy and Takeuchi (2022). Each polyhedron in the $n$-dimensional space corresponds to an interval in the one-dimensional space $\mathcal{Z}$. Thus, we can examine these intervals sequentially in $\mathcal{Z}$ to determine whether they yield the same selected anomalous regions as observed. To this end, we need to identify a subinterval $\mathcal{Z}^{\mathrm{sub}}\subseteq\mathcal{Z}$. We show that the anomalous region from a diffusion model can be characterized by a piecewise-linear mapping, and that for each $z\in\mathbb{R}$, the subinterval $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ can be computed analytically by solving a system of linear inequalities (see Appendix B for details).
4.3 Identification of $\mathcal{Z}$